# Disorder-Engineered Hybrid Plasmonic Cavities for Emission Control of Defects in hBN

### v1

@@ Abstract
Defect-based quantum emitters in hexagonal boron nitride (hBN) are promising building blocks for scalable quantum photonics due to their stable single-photon emission at room temperature. However, enhancing their emission intensity and controlling the decay dynamics remain significant challenges. This study demonstrates a low-cost, scalable fabrication approach to integrate plasmonic nanocavities with defect-based quantum emitters in hBN nanoflakes. Using the thermal dewetting process, we realize two distinct configurations: stochastic Ag nanoparticles (AgNPs) on hBN flakes and hybrid plasmonic nanocavities formed by AgNPs on top of hBN flakes supported on gold/silicon dioxide ($\rm Au/SiO_{2})$ substrates. While AgNPs on bare hBN yield up to a two-fold photoluminescence (PL) enhancement with reduced emitter lifetimes, the hybrid nanocavity architecture provides a dramatic, up to 100-fold PL enhancement and improved uniformity across multiple emitters, all without requiring deterministic positioning. Finite-difference time-domain (FDTD) simulations and time-resolved PL measurements confirm size-dependent control over decay dynamics and cavity-emitter interactions. Our versatile solution overcomes key quantum photonic device development challenges, including material integration, emission intensity optimization, and spectral multiplexity. Future work will explore potential applications in integrated photonic circuits hosting on-chip quantum systems and hBN-based label-free single-molecule detection through such quantum nanoantennas.
Keywords: hBN, photoluminescence, nanocavity, enhancement, dewetting.
I Introduction
@@
Quantum emitters are central to quantum technologies due to their intrinsically non-classical light emission [1]. Emitters based on two-dimensional (2D) materials offer a promising platform for scalable photonic quantum technologies, owing to their compactness, stability, and tunable properties [2]. Among these materials, hexagonal boron nitride (hBN) offers several advantageous properties, such as high chemical and thermal stability [3, 4], a wide bandgap (6 eV), and the ability to host room-temperature single-photon emitters based on atomic defects [5, 6, 7, 8]. Despite some challenges, such as defect control, spectral variability, and quantum yield limitations, advances in material synthesis, defect engineering, and nanofabrication are progressively overcoming these hurdles, paving the way for the integration of hBN into next-generation quantum photonic devices [9, 10, 11]. Overall, these properties also make hBN an attractive material for cavity systems. Its wide bandgap effectively isolates quantum emitters from environmental noise, enhancing emission stability, while its ultra-thin, atomically smooth layers enable precise control over cavity geometries [12, 13]. However, similar to the other 2D materials, the main challenges of using hBN as an emitter in a cavity system are the precise defect placement and uniform layer thickness, which remain significant technical hurdles [14, 15]. Additionally, integrating hBN with metallic or dielectric substrates requires advanced fabrication techniques to preserve its intrinsic properties, adding complexity to the manufacturing process.
@@
The idea of coupling a quantum emitter with a photonic cavity originates from modifying spontaneous emission by changing the environment of the quantum emitter (Purcell effect [16]), which can be used to adjust its emission characteristics. After the first experimental demonstration of modified emission from a single quantum system with a millimeter-sized cavity and with the advances in microfabrication techniques, microcavities came forth due to their comparatively diminutive volumes in the micron range, with the potential to integrate into large-scale photonic circuitry [17].
@@
Enhancement Factor Precise Location of hBN to Cavity/Nanostructure REF
@@
11 Emitters are located at the center of a reflective dielectric cavity for optimized reflection. [18]
@@
3 hBN emitters positioned near gold (Au) nanospheres using atomic force microscopy. [19]
@@
1.4–4.0 Emitters transferred from dielectric to metallic substrates to form a hybrid metal-dielectric structure. [20]
@@
250 Emitters are placed in a nano-patch antenna cavity coupled tightly to the cavity area. [21]
8 Carbon enriched hBN layers (hBN:C) were transferred from the hBN:C crystal to a silicon dioxide/silicon (${\rm SiO}_{2}$/Si) substrate. [22]
@@
2 Solution-synthesized Ag nanocubes are drop-casted on ion-implanted ultrathin hBN flakes on Si substrate. [23]
@@
Metallic nanostructures featuring nanoscale gaps support hybrid plasmonic modes with ultra-small mode volumes and highly localized electromagnetic fields, making them ideal for exploring light-matter interactions at the nanoscale [24, 25]. Despite their potential, conventional cavity designs, such as gaps of plasmonic bowtie antennas and dielectric resonators, encounter significant constraints in practical applications. While bowtie antennas are efficient in field concentration, their fabrication is challenging due to the minuscule gap, and they often lack scalability, making them unsuitable for widespread use [26, 27]. Similarly, dielectric resonators exhibit higher mode volumes and provide less field localization and limited Purcell factor enhancement [28].
@@
Additionally, both plasmonic and dielectric systems often struggle with integration into 2D material systems due to mismatched geometries, reduced stability, and inconsistent electromagnetic coupling efficiencies [29, 30]. Alternative antenna designs, including dipole, monopole, and Yagi-Uda antennas, address some limitations but introduce their challenges. Dipole and monopole antennas, while simple and efficient, cannot localize fields at nanoscale volumes compared to plasmonic designs, limiting their effectiveness in enhancing light-matter interactions [31]. Although Yagi-Uda antennas offer excellent directional emission and gain, they are challenging to miniaturize for nanoscale systems, complicating their integration with 2D material-based quantum technologies [32, 33]. Furthermore, their alignment and coupling requirements demand precise fabrication, increasing overall complexity and cost.
@@
Table 1 lists some of the plasmonic enhancement factors of hBN PL as reported in the literature. In studies concerning cavity-enhanced emission from hBN quantum emitters, enhancement factors (EF) exhibit considerable variation based on the type of cavity and the location of the emitter. Dielectric cavities often produce moderate electric field (EF) values (e.g., 11 in [18]), but plasmonic structures like nano-patch antennas can attain significantly greater enhancements (up to 250 in [21]) owing to intense field confinement and near-field coupling. Conversely, systems dependent on random nanoparticle distribution or weak hybrid interfaces frequently yield modest enhancement factors (EF < 4), highlighting the significance of accurate emitter-cavity alignment and resonant mode optimization. Similarly, dielectric optical cavities have also been used to enhance the PL of hBN [34, 35]. The recorded enhancement factors by dielectric cavities are comparable to those of the plasmonic nanocavities. One promising technique for integrating emitters with plasmonic nanostructures is the thermal dewetting of noble metal films. In this process, a thin metal film breaks up into nano-islands upon annealing, forming stochastic nanoscale antennas near the emitter locations without requiring lithography or precise placement. These disorder-engineered, self-assembled nano-antennas exhibit coupling to nearby defect emission in hBN flakes with a minimal fabrication effort.
In this study, we explore two complementary approaches to control the emission from single defects in hBN using plasmonic cavities: (i) direct formation of AgNPs via thermal dewetting of a Ag film on hBN flakes dispersed on a Si substrate and (ii) fabrication of hybrid plasmonic nanocavities by thermal dewetting of a Ag film on hBN flakes dispersed on $\rm Au/SiO_{2}$ coated Si substrates. We investigate the impact of these plasmonic configurations on key emission properties such as photoluminescence intensity, lifetime, saturation behavior, and single-photon purity through a combination of experimental measurements and FDTD simulations.
II Results and Discussion
Figure 1a shows a typical optical image of the sample surface, where several hBN structures are clearly visible. A scanning electron microscope (SEM) image of the selected bulk hBN flakes is shown in Figure 1b. The hBN sample consists of multiple flakes with varying sizes (on the order of tens of nanometers) and orientations, which may significantly influence the optical properties of the quantum emitters present within the flakes. To investigate the spectral characteristics of these structures, a home-built micro-photoluminescence ($\mu$-PL) setup was employed. The details of the setup can be found in the Experimental Section.
@@ II Results and Discussion
Figure 1c shows a representative room-temperature PL spectrum of a bulk hBN flake excited with a 532 nm continuous-wave (CW) laser. The spectrum exhibits several characteristic features: a Raman peak at 547 nm corresponding to the Si substrate (520 ${\rm cm^{-1}}$), and a sharp Raman peak at 574 nm (1370 ${\rm cm^{-1}}$) arising from the high-energy $E_{2g}$ phonon mode of hBN. Although hBN also possesses a low-energy Raman-active mode around 52 ${\rm cm^{-1}}$, this feature is not visible in our measurements due to strong spectral filtering near the excitation wavelength [36, 37, 38]. A sharp emission peak centered at 684 nm, accompanied by a weaker and broader feature at 753 nm, is also observed. These are assigned to the zero-phonon line (ZPL) and its corresponding phonon sideband (PSB), respectively, originating from a single defect in hBN. The energy difference between the ZPL and PSB matches well with the energy of the high-energy phonon mode, confirming their assignment. Polarization-resolved measurements (inset of Fig. 1c) reveal strong linear polarization in both excitation (filled circles) and emission (open circles), consistent with a single dipole transition [39, 40, 41].
@@
To confirm the quantum nature of the emission, we measured the second-order photon correlation function $g^{(2)}(\tau)$ using a Hanbury-Brown and Twiss (HBT) setup comprising a 50:50 beamsplitter and two single-photon counting modules (SPCMs), with timing recorded via a multichannel time-stamping unit. Figure 1d shows the histogram of detection time delays under CW excitation. The pronounced antibunching dip at zero delay with $g^{(2)}(0)=0.3$ confirms the single-photon nature of the emission. The observed bunching behavior indicates a three-level system, a feature commonly observed in solid-state quantum emitters, including defects in hBN [39] and diamond [42]. A schematic of the emitter’s level structure is shown in the inset of Fig. 1c. It includes a ground state $\lvert 1\rangle$, an excited state $\lvert 2\rangle$, and a metastable shelving state $\lvert 3\rangle$. The relevant transition rates are denoted as follows: $\gamma_{12}$ (excitation), $\gamma_{21}$ (radiative $\gamma_{r}$ and nonradiative $\gamma_{nr}$ decay), $\gamma_{23}$ (shelving), and $\gamma_{31}$ (deshelving). These rates govern the dynamics and intensity of the ZPL emission. The deviation of $g^{(2)}(0)$ from the ideal value of zero is attributed to imperfect spectral filtering of the ZPL and the limited temporal resolution of the HBT system (approximately 400 ps), which is insufficient to fully resolve fast decay processes governed by $\gamma_{21}$ and $\gamma_{12}$. To mitigate these limitations, we also performed photon correlation measurements under pulsed excitation at 483 nm. Under these conditions, a much stronger suppression of the peak at the zero time-delay is observed, with an antibunching value of $g^{(2)}(0)=0.1$, as shown in the inset of Fig. 1d).
@@
In order to investigate the effect of Ag nanoantennas on the optical properties of single defects in hBN, the nanoantennas were fabricated in the vicinity of the defects using a self-organized approach, rather than moving the defect centers close to pre-fabricated antennas [19]. For this, a technique called solid-state dewetting of thin films was employed. A 5/13 nm thick Ag film was deposited on the hBN-coated silicon substrate and subsequently annealed at approximately $350\,^{\circ}$[43, 44, 45]. Silver was chosen due to its low optical loss, strong and narrow plasmonic resonance bands, and superior dewetting properties compared to other noble metals such as gold. By controlling the dewetting duration, the size distribution of the resulting nanoantennas can be tuned (Supporting Information Figure S1).
During the annealing process, the thin Ag film becomes unstable at temperatures well below the melting point of silver due to a high surface-area-to-volume ratio [46]. To minimize this instability, silver atoms diffuse and agglomerate, leading to the formation of nano-sized islands as presented in Figure 2a. These nanoislands take on a largely hemispheroidal shape, resulting in Ag nanoantennas distributed across the hBN flakes as well as on the silicon substrate. Because the dewetting process is stochastic, nanoantennas also form in close proximity to defect centers.
The fabrication process consists of two steps: vacuum thermal metal evaporation followed by thermal annealing in a nitrogen environment. This method is gentle on both the hBN flakes and the silicon substrate, which are stable up to $\sim 800\,^{\circ}$[47]. By controlling the initial Ag film thickness, the size distribution of the resulting nanoantennas can be tuned. In this work, 5 nm and 13 nm thick films produced nanoantennas with average diameters of approximately 35 nm and 110 nm, respectively (Figure 2b and c). These size regimes were selected to demonstrate both quenching and enhancement effects on the fluorescence of single defects in hBN, as the size of Ag nanoantennas strongly influences their absorption and scattering properties [48].
II.1 AgNPs on hBN: Emitter Characterization and Plasmonic Coupling
@@
Figure 3: (a) PL spectra of a single defect in hBN with a ZPL at 663 nm, measured before and after the deposition of Ag nanoantennas. PSBs appear around 730 nm, while the peak at 574 nm corresponds to Raman scattering of hBN. A significant quenching of ZPL intensity is observed after Ag deposition. (b) Excitation power-dependent intensities of the ZPL and hBN Raman scattering before and after Ag nanoantenna formation. ZPL emission shows strong quenching, while the Raman peak is enhanced due to near-field enhancement effects. (c) Time-resolved PL measurements showing a clear reduction in lifetime after Ag nanoantenna formation, indicating enhanced total decay rate due to coupling with plasmonic modes. (d) FDTD simulation results for a model system consisting of a defect in hBN coupled to a 40 nm diameter Ag nanoparticle. Radiative and non-radiative decay rates are plotted as functions of emitter–nanoparticle distance. Radiative enhancement dominates at intermediate distances, while non-radiative decay sharply increases at shorter separations. Higher modification is observed for dipoles oriented parallel to the Ag surface, while lower rates occur for orthogonal dipole orientations as shown in the shaded regions.
@@
To investigate the origin of the observed quenching in PL emission under identical excitation conditions, power-dependent PL measurements were conducted on the same defect emitting at 663 nm, both before and after AgNP fabrication. From each spectrum, the intensity of the ZPL emission and the hBN Raman scattering were extracted and are shown in Figure 3b for quantitative comparison. While the Raman signal is visible only at high excitation powers, it exhibits a linear power dependence, consistent with its scattering origin. Notably, a clear enhancement in the Raman scattering is observed after the AgNP formation, attributed to surface-enhanced Raman scattering (SERS) [49, 50, 51]. In contrast, the ZPL intensity shows typical saturation behavior. At low excitation powers, the ZPL intensity $I$ increases linearly with power, following $I\approx\gamma_{12}\times QE$, where $\gamma_{12}$ is the excitation rate and $QE$ is the quantum efficiency. At higher excitation powers, saturation sets in, and the intensity reaches a maximum limited by the radiative and non-radiative decay rates, as well as the involvement of metastable triplet states. In this regime, the PL intensity can be described by $I=\gamma_{r}\gamma_{31}/(\gamma_{31}+\gamma_{23})$, where $\gamma_{r}$ is the radiative decay rate, and $\gamma_{23},\gamma_{31}$ are the intersystem crossing rates associated with the triplet states. The quantum efficiency, defined as $QE=\gamma_{r}/(\gamma_{nr}+\gamma_{r}+\gamma_{23})$ reflects the competition between radiative decay ($\gamma_{r}$), non-radiative decay ($\gamma_{nr}$), and intersystem crossing. The significant quenching observed after AgNP fabrication is consistent with an increase in non-radiative decay pathways, likely caused by energy transfer from the emitter to the nearby metallic nanoparticle, which reduces the $QE$.

### v2

@@ Abstract
Defect-based quantum emitters in hexagonal boron nitride (hBN) are promising building blocks for scalable quantum photonics due to their stable single-photon emission at room temperature. However, enhancing their emission intensity and controlling the decay dynamics remain significant challenges. This study demonstrates a low-cost, scalable fabrication approach to integrate plasmonic nanocavities with defect-based quantum emitters in hBN nanoflakes. Using the thermal dewetting process, we realize two distinct configurations: stochastic Ag nanoparticles (AgNPs) on hBN flakes and hybrid plasmonic nanocavities formed by AgNPs on top of hBN flakes supported on gold/silicon dioxide (Au/SiO2) substrates. While AgNPs on bare hBN yield up to a two-fold photoluminescence (PL) enhancement with reduced emitter lifetimes, the hybrid nanocavity architecture provides a dramatic, up to 100-fold PL enhancement and improved uniformity across multiple emitters, all without requiring deterministic positioning. Finite-difference time-domain (FDTD) simulations and time-resolved PL measurements confirm size-dependent control over decay dynamics and cavity-emitter interactions. Our versatile solution overcomes key quantum photonic device development challenges, including material integration, emission intensity optimization, and spectral multiplexity.
[…]
Keywords: single photon source, photoluminescence, nanocavity, dewetting
1 Introduction
@@
Quantum emitters are central to quantum technologies due to their intrinsically non-classical light emission Fox 2025. Emitters based on two-dimensional (2D) materials offer a promising platform for scalable photonic quantum technologies, owing to their compactness, stability, and tunable properties Kianinia et al. 2022. Among these materials, hexagonal boron nitride (hBN) offers several advantageous properties, such as high chemical and thermal stability Grosso et al. 2017, Herzig Sheinfux et al. 2024, a wide bandgap (6 eV), and the ability to host room-temperature single-photon emitters based on atomic defects Yamamura et al. 2024, Prasad et al. 2024, Tran et al. 2016, Çakan et al. 2025. Despite some challenges, such as defect control, spectral variability, and quantum yield limitations, advances in material synthesis, defect engineering, and nanofabrication are progressively overcoming these hurdles, paving the way for the integration of hBN into next-generation quantum photonic devices Hennessey et al. 2025, Nonahal et al. , Li et al. 2021. Overall, these properties also make hBN an attractive material for cavity systems. Its wide bandgap effectively isolates quantum emitters from environmental noise, enhancing emission stability, while its ultra-thin, atomically smooth layers enable precise control over cavity geometries Kim et al. 2018, Shandilya et al. 2019. However, similar to the other 2D materials, the main challenges of using hBN as an emitter in a cavity system are the precise defect placement and uniform layer thickness, which remain significant technical hurdles Tran et al. 2017, Mendelson et al. 2019. Additionally, integrating hBN with metallic or dielectric substrates requires advanced fabrication techniques to preserve its intrinsic properties, adding complexity to the manufacturing process.
@@
The idea of coupling a quantum emitter with a photonic cavity originates from modifying spontaneous emission by changing the environment of the quantum emitter (Purcell effect Purcell 1995), which can be used to adjust its emission characteristics. After the first experimental demonstration of modified emission from a single quantum system with a millimeter-sized cavity and with the advances in microfabrication techniques, microcavities came forth due to their comparatively diminutive volumes in the micron range, with the potential to integrate into large-scale photonic circuitry Megahd et al. 2022.
@@
Enhancement Factor Precise Location of hBN to Cavity/Nanostructure Ref.
@@
11 Emitters are located at the center of a reflective dielectric cavity for optimized reflection. Zeng et al. 2023
@@
3 hBN emitters positioned near gold (Au) nanospheres using atomic force microscopy. Nguyen et al. 2018
@@
1.4–4.0 Emitters transferred from dielectric to metallic substrates to form a hybrid metal–dielectric structure. Gérard et al.
$\sim$4 Emitters localized ($\sim$80 nm) in exfoliated few-layer hBN flakes on lacey carbon grids (relative to NV centers, extremely bright UV SPE at 4.1 eV). Bourrellier et al. 2016
<1<1 Emitters located in suspended single-crystal hBN membranes. Exarhos et al. 2017
$\sim$10 Carbon incorporation (MOVPE, MBE, HOPG conversion, ion implantation) tunes PL intensity. Mendelson et al. 2021
@@
250 Emitters are placed in a nano-patch antenna cavity tightly coupled to the cavity region. Xu et al. 2022
8 Carbon-enriched hBN layers (hBN:C) transferred onto a SiO2/Si substrate. Karanikolas et al. 2023
@@
2 Solution-synthesized Ag nanocubes drop-casted on ion-implanted ultrathin hBN flakes on Si substrate. Dowran et al. 2023
@@
Metallic nanostructures featuring nanoscale gaps support hybrid plasmonic modes with ultra-small mode volumes and highly localized electromagnetic fields, making them ideal for exploring light-matter interactions at the nanoscale Barreda et al. 2021, Zhang et al. . Despite their potential, conventional cavity designs, such as gaps of plasmonic bowtie antennas and dielectric resonators, encounter significant constraints in practical applications. While bowtie antennas are efficient in field concentration, their fabrication is challenging due to the minuscule gap, and they often lack scalability, making them unsuitable for widespread use Hu et al. 2018, Khalil et al. 2021. Similarly, dielectric resonators exhibit higher mode volumes and provide less field localization and limited Purcell factor enhancement Dmitriev et al. 2023.
@@
Additionally, both plasmonic and dielectric systems often struggle with integration into 2D material systems due to mismatched geometries, reduced stability, and inconsistent electromagnetic coupling efficiencies Meng et al. , Yan et al. 2020. Alternative antenna designs, including dipole, monopole, and Yagi-Uda antennas, address some limitations but introduce their challenges. Dipole and monopole antennas, while simple and efficient, cannot localize fields at nanoscale volumes compared to plasmonic designs, limiting their effectiveness in enhancing light-matter interactions Patel and Argyropoulos 2015. Although Yagi-Uda antennas offer excellent directional emission and gain, they are challenging to miniaturize for nanoscale systems, complicating their integration with 2D material-based quantum technologies Maksymov et al. 2012, Coenen et al. 2014. Furthermore, their alignment and coupling requirements demand precise fabrication, increasing overall complexity and cost.
@@
Table 1 lists some of the plasmonic enhancement factors of hBN PL as reported in the literature. In studies concerning cavity-enhanced emission from hBN quantum emitters, enhancement factors (EF) exhibit considerable variation based on the type of cavity and the location of the emitter. Dielectric cavities often produce moderate enhancement factor values (e.g., 11 in Zeng et al. 2023), but plasmonic structures like nano-patch antennas can attain significantly greater enhancements (up to 250 in Xu et al. 2022) owing to intense field confinement and near-field coupling. Conversely, systems dependent on random nanoparticle distribution or weak hybrid interfaces frequently yield modest enhancement factors (EF <4<4), highlighting the significance of accurate emitter-cavity alignment and resonant mode optimization. Similarly, dielectric optical cavities have also been used to enhance the PL of hBN fröch, Häußler et al. 2021. The recorded enhancement factors by dielectric cavities are comparable to those of the plasmonic nanocavities. One promising technique for integrating emitters with plasmonic nanostructures is the thermal dewetting of noble metal films. In this process, a thin metal film breaks up into nano-islands upon annealing, forming stochastic nanoscale antennas near the emitter locations without requiring lithography or precise placement. These disorder-engineered, self-assembled nano-antennas exhibit coupling to nearby defect emission in hBN flakes with a minimal fabrication effort.
In this study, multilayer hBN flakes dispersed in solution were used, drop-casted onto the target substrates. This choice was motivated by the ease of integration, scalability, and compatibility with various substrate types, in contrast to exfoliated or CVD-grown flakes, which often require more elaborate transfer or growth protocols. Drop-casting enables rapid prototyping of emitter–cavity configurations without complex positioning, aligning with our goal of exploring lithography-free plasmonic architectures for quantum photonic applications. While this approach sacrifices some control over flake thickness and orientation, it provides a pragmatic pathway towards device concepts that can be implemented in less resource-intensive fabrication environments. Therefore, we explore two complementary approaches to control the emission from single defects in hBN using plasmonic cavities: (i) direct formation of AgNPs via thermal dewetting of a Ag film on hBN flakes dispersed on a Si substrate and (ii) fabrication of hybrid plasmonic nanocavities by thermal dewetting of a Ag film on hBN flakes dispersed on $\rm Au/SiO_{2}$ coated Si substrates. We investigate the impact of these plasmonic configurations on key emission properties such as photoluminescence intensity, lifetime, saturation behavior, and single-photon purity through a combination of experimental measurements and FDTD simulations. We remark that identifying the exact microscopic defect configuration responsible for each emission line in hBN is beyond the scope of this study. Given the well-known heterogeneity of hBN emitters, our focus is on experimentally demonstrating scalable plasmonic coupling and decay-dynamics control, supported by numerical simulations, rather than assigning a unique atomic defect species to each emitter.
2 Results and discussion
[…]
It is important to note that the color centers studied here correspond to naturally occurring intrinsic point defects in hBN, whose fundamental optical characteristics are not dictated by the flake preparation method. Previous reports have demonstrated comparable single-photon emission properties in exfoliated, CVD-grown, and solution-processed hBN flakes, including similar zero-phonon line features, polarization behavior, and photon antibunching Kianinia et al. 2022, Tran et al. 2016, Mendelson et al. 2019. Accordingly, the focus of the present study is demonstrating scalable plasmonic coupling and decay-dynamics control, supported by numerical simulations, rather than deterministic control of flake thickness or defect identity.
Figure 1a shows a typical optical image of the sample surface, where several hBN structures are clearly visible. Figure 1b shows a scanning electron microscope (SEM) image of the selected bulk hBN flakes, revealing that the drop-casting process results in heterogeneous morphologies, including stacked multilayer regions and significant variations in lateral size and thickness. While such morphology differs from the atomically flat and thickness-controlled flakes obtainable via exfoliation or CVD growth, it reflects a realistic and scalable material platform. While exfoliated or CVD-grown hBN flakes indeed offer superior thickness control and atomically flat interfaces, the present work deliberately employs solution-processed flakes to evaluate plasmonic coupling under fabrication-tolerant and scalable conditions, where perfect 2D interfaces cannot be assumed. In the present work, this morphological variability is not treated as a limitation but serves to examine plasmonic coupling in non-ideal and heterogeneous material environments. To investigate the spectral characteristics of these structures, a home-built micro-photoluminescence ($\mu$-PL) setup was employed. The details of the setup can be found in the Experimental Section.
@@ 2 Results and discussion
Figure 1c shows a representative room-temperature PL spectrum of a bulk hBN flake excited with a 532 nm continuous-wave (CW) laser. The spectrum exhibits several characteristic features: a Raman peak at 547 nm corresponding to the Si substrate (520 $\rm cm^{-1}$}), and a sharp Raman peak at 574 nm (1370 $\rm cm^{-1}$}) arising from the high-energy $E_{2g}$ phonon mode of hBN. Although hBN also possesses a low-energy Raman-active mode around 52 $\rm cm^{-1}$}, this feature is not visible in our measurements due to strong spectral filtering near the excitation wavelength Kuzuba et al. 1978, Nemanich et al. 1981, Arı et al. 2025. A sharp emission peak centered at 684 nm, accompanied by a weaker and broader feature at 753 nm, is also observed. These features are assigned to the zero‑phonon line (ZPL)—a sharp emission arising from a phonon‑less electronic transition—and its corresponding phonon sideband (PSB), which originates from emission involving phonon-assisted relaxation. This assignment is consistent with prior observations in hexagonal boron nitride, where the ZPL appears as a narrow peak while the PSB manifests as a broader spectral feature Arı et al. 2025, Denning et al. 2020. The energy difference between the ZPL and PSB matches well with the energy of the high-energy phonon mode, confirming their assignment. Polarization-resolved measurements (inset of Fig. 1c) reveal strong linear polarization in both excitation (filled circles) and emission (open circles), consistent with a single dipole transition Tran et al. 2015, Jungwirth and Fuchs 2017, Kumar et al. 2024.
To confirm the quantum nature of the emission, the second-order photon correlation function $g^{(2)}(\tau)$ was measured using a Hanbury Brown and Twiss (HBT) interferometer consisting of a 50:50 beamsplitter and two single-photon counting modules (SPCMs), with timing recorded via a multichannel time- stamping unit. Here, a “single defect” refers to an optically isolated individual emitter within the confocal excitation and collection volume, rather than a deterministically created or structurally isolated atomic defect. Figure 1d shows the photon correlation histogram recorded under continuous-wave (CW) excitation at room temperature. The pronounced antibunching dip at zero time delay, with $g^{(2)}(0)=0.3$, confirms sub-Poissonian photon statistics and thus single-photon emission from an individual emitter.
@@
The bunching behavior observed at longer delay times is characteristic of a three-level system, which is commonly reported for solid-state quantum emitters, including defects in hBN Tran et al. 2015 and diamond Kurtsiefer et al. 2000. A schematic representation of the corresponding energy-level structure is shown in the inset of Fig. 1c, comprising a ground state $\lvert 1\rangle$, an excited state $\lvert 2\rangle$, and a metastable shelving state $\lvert 3\rangle$. The relevant transition rates are denoted as follows: $\gamma_{12}$ (excitation), $\gamma_{21}$ (radiative $\gamma_{r}$ and nonradiative $\gamma_{nr}$ decay), $\gamma_{23}$ (shelving), and $\gamma_{31}$ (deshelving). These rates govern the dynamics and intensity of the ZPL emission. The observed deviation of $g^{(2)}(0)$ from the ideal value of zero is attributed to imperfect spectral filtering of the zero-phonon line and the finite temporal resolution of the HBT setup (approximately 400 ps), which limits the ability to fully resolve fast decay processes Tran et al. 2015. To further validate the single-photon character of the emission, photon correlation measurements were also performed under pulsed excitation at 483 nm. Under these conditions, a stronger suppression of the zero-delay peak is observed, yielding an antibunching value of $g^{(2)}(0)=0.1$, as shown in the inset of Fig. 1d.
@@
AgNPs on hBN: Emitter Characterization and Plasmonic Coupling
The nanoparticle sizes investigated here ($\sim$ 35 nm and $\sim$ 110 nm) were chosen to explore two distinct plasmonic regimes. Silver was selected as the plasmonic material due to its superior optical performance in the visible spectral range, where hBN defect emission occurs. Among noble metals, Ag exhibits lower intrinsic optical losses and supports sharper localized surface plasmon resonances compared to Au, enabling stronger near-field enhancement and more efficient modification of radiative decay rates. In addition, Ag films exhibit favorable solid-state dewetting behavior at relatively low annealing temperatures, allowing lithography-free and scalable formation of plasmonic nanoantennas with controlled size distributions. Particles smaller than $\sim$ 30 nm, while supporting strong local fields, typically have lower scattering cross-sections and higher ohmic losses Jain et al. 2007, which can limit the radiative contribution to PL enhancement. On the other hand, particles significantly larger than $\sim$ 100 nm can exhibit multipolar plasmon modes and increased radiative damping Kelly et al. 2003, reducing spectral overlap with the emitter’s ZPL and altering cavity mode profiles. Optimization across a broader size range may further tailor enhancement for specific applications.
@@
Figure 3: (a) PL spectra of a single defect in hBN with a ZPL at 663 nm, measured before and after the deposition of Ag nanoantennas. PSBs appear around 730 nm, while the peak at 574 nm corresponds to Raman scattering of hBN. A significant quenching of ZPL intensity is observed after Ag deposition. (b) Excitation power-dependent intensities of the ZPL and hBN Raman scattering before and after Ag nanoantenna formation. ZPL emission shows strong quenching (are shown with solid circles and squares), while the Raman peak is enhanced due to near-field enhancement effects (are indicated by open circles and squares). (c) Time-resolved PL measurements showing a clear reduction in lifetime after Ag nanoantenna formation, indicating enhanced total decay rate due to coupling with plasmonic modes. (d) FDTD simulation results for a model system consisting of a defect in hBN coupled to a 40 nm diameter AgNP. Radiative and non-radiative decay rates are plotted as functions of emitter–nanoparticle distance. Radiative enhancement dominates at intermediate distances, while non-radiative decay sharply increases at shorter separations. Higher modification is observed for dipoles oriented parallel to the Ag surface, while lower rates occur for orthogonal dipole orientations as shown in the shaded regions.
@@
To investigate the origin of the observed quenching in PL emission under identical excitation conditions, power-dependent PL measurements were conducted on the same defect emitting at 663 nm, both before and after AgNP fabrication. From each spectrum, the intensity of the ZPL emission and the hBN Raman scattering were extracted and are shown in Figure 3b for quantitative comparison. While the Raman signal is visible only at high excitation powers, it exhibits a linear power dependence, consistent with its scattering origin. Notably, a clear enhancement in the Raman scattering is observed after the AgNP formation, attributed to surface-enhanced Raman scattering (SERS) Fleischmann et al. 1974, Jeanmaire and Van Duyne 1977, Kneipp et al. 1997. In contrast, the ZPL intensity shows typical saturation behavior. At low excitation powers, the ZPL intensity $I$ increases linearly with power, following $I\approx\gamma_{12}\times QE$, where $\gamma_{12}$ is the excitation rate and $QE$ is the quantum efficiency Kurtsiefer et al. 2000, Aharonovich et al. 2016. At higher excitation powers, saturation sets in, and the intensity reaches a maximum limited by the radiative and non-radiative decay rates, as well as the involvement of metastable triplet states. In this regime, the PL intensity can be described by $I=\gamma_{r}\gamma_{31}/(\gamma_{31}+\gamma_{23})$, where $\gamma_{r}$ is the radiative decay rate, and $\gamma_{23},\gamma_{31}$ are the intersystem crossing rates associated with the triplet states. The quantum efficiency, defined as $QE=\gamma_{r}/(\gamma_{nr}+\gamma_{r}+\gamma_{23})$ reflects the competition between radiative decay ($\gamma_{r}$), non-radiative decay ($\gamma_{nr}$), and intersystem crossing. The significant quenching observed after AgNP fabrication is consistent with an increase in non-radiative decay pathways, likely caused by energy transfer from the emitter to the nearby metallic nanoparticle, which reduces the $QE$.